OpenStudy (anonymous):

how do i factor x^a+2-2x^a+1-3x^a?

OpenStudy (anonymous):

\[x^{a} + 2 - 2x^{a}+1-3x^{a}\]Is this your question?

OpenStudy (raden):

or x^(a+2)-2x^(a+1)-3x^a ?

OpenStudy (anonymous):

First arrange the like terms.i.e.\[x^{a}-2x^{a}-3x^{a}+2+1\] \[x^{a}-2x^{a}-3x^{a} = -4x^{a}\]\[x^{a}-2x^{a}-3x^{a}+2+1 = -4x^{a}+3 = 3 - 4x^{a}\]

OpenStudy (anonymous):

thnks

OpenStudy (anonymous):

you are welcomed...